what are all ordered pairs of numbers (x, y) which satisfy x^2 – xy + y^2 = 7 and x- xy + y = -1 ?

Answer by richard1234(7193)

Add 3xy to both sides of the 1st equation.



The second equation is equivalent to

. Substituting this into the first equation,





This is a quadratic in x+y, and the solutions are x+y = 5 and x+y = -2. If x+y = 5, substituting into the second equation we obtain



And if x+y = -2, substituting into the second equation yields



Therefore we have two cases: x+y = 5 & xy = 6, and x+y = -2 & xy = -1. The first case yields {3,2}; the second case we can let x = -2-y, in which



By the quadratic formula, , the solution obtained is